### v1

# TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremenal Learning

Abstract
Continual Learning is an unresolved challenge, whose relevance increases when considering modern applications. Unlike the human brain, trained deep neural networks suffer from a phenomenon called Catastrophic Forgetting, where they progressively lose previously acquired knowledge upon learning new tasks. To mitigate this problem, numerous methods have been developed, many relying on replaying past exemplars during new task training. However, as the memory allocated for replay decreases, the effectiveness of these approaches diminishes. On the other hand, maintaining a large memory for the purpose of replay is inefficient and often impractical. Here we introduce TEAL, a novel approach to populate the memory with exemplars, that can be integrated with various experience-replay methods and significantly enhance their performance on small memory buffers. We show that TEAL improves the average accuracy of the SOTA method XDER as well as ER and ER-ACE on several image recognition benchmarks, with a small memory buffer of 1-3 exemplars per class in the final task. This confirms the hypothesis that when memory is scarce, it is best to prioritize the most typical data.
1 Introduction
With the recent advances in deep neural networks, research interest in incremental learning has increased. Here, the problem of catastrophic forgetting, as defined by McCloskey and Cohen [20], can be severe. The necessity to incorporate new task knowledge into an already trained network has become more relevant given that training over extensive datasets is time-consuming. It is often impractical to retrain the network from scratch on the combined original and new task data. Furthermore, access to the already-trained data might be limited or non-existent.
Figure 1: Catastrophic forgetting when training CIFAR-10 [13] sequentially with only finetuning.
Various methods have been developed to tackle this challenge across different frameworks, categorized by Van de Ven et al. [25] into three types: task-incremental, domain-incremental, and class-incremental learning. The core concept of incremental learning is that a model must learn tasks sequentially, one after another. Among these, class-incremental learning is considered the most challenging. In this scenario, each task introduces new classes, and the model must identify the specific class of each input without access to the task ID on which it was trained.
There are several approaches to incremental learning, each with its own distinct assumptions and configurations. In this paper, we adhere to the constrained setup described by De Lange et al. [7], which does not rely on task boundaries during training or testing. This setup maintains a constant memory size throughout incremental training, ensuring that it does not exceed a predetermined limit established at the outset. At each incremental step, we can add new exemplars to the memory buffer only after space has been allocated for them.
We focus our attention on a prevalent and rather successful framework called Experience Replay (ER), which involves storing a set of exemplars in memory and reusing them for rehearsal purposes while training on new tasks. Within this framework, several strategies exist for selecting which exemplars to retain in memory. Masana et al. [19] demonstrate that the most successful strategies are either random-sampling or Herding as defined by Welling [27]. Not surprisingly, the smaller the memory buffer is, the less effective the strategy is at mitigating catastrophic forgetting. This leads us to the following question: Are these strategies necessarily optimal for all memory sizes? In other words, could different strategies prove to be more effective depending on the size of the memory buffer?
Our proposed method, TEAL, operates on similar principles, aiming to identify a set of representative exemplars that also exhibit diversity. An exemplar is deemed representative if its likelihood, when considered within the distribution of all points in an appropriate latent space, is high. Leveraging clustering ensures the selection of a diverse set of exemplars. This approach enhances the model’s capacity to retain previous knowledge while effectively learning new tasks, even when constrained by a small memory buffer. Central to the success of this approach is our ability to derive an appropriate latent space. The subsequent sections elaborate on our methodology, experimental setup, comparative performance against alternative selection strategies, and the enhancements TEAL offers to various existing class-incremental methods.
Related work    Several approaches exist for incremental learning [see 7]. Alongside ER, alternatives include Generative Replay (GR) [23], Parameter Isolation, and Regularization. Similar to ER, GR replays data from previous tasks during new task training, but uses a generative model to create new samples instead of retaining exemplars seen by the model [6, 8, 9]. Parameter Isolation assigns distinct parameters to each task to reduce forgetting [18] by fixing parameters assigned to previous tasks, while Regularization-based methods [15] incorporate additional terms into the loss function to retain prior knowledge while learning from new data.
These methods differ in how they utilize memory: ER stores exemplars, GR stores a generative model, and Parameter Isolation stores task-specific parameters. Regularization-based methods do not rely on memory. This diversity makes it challenging to compare methods directly. In particular, we note that since generative models tend to be very large, this approach is hardly suitable to the domain of small memory buffers addressed below. Hence, our focus in this paper is on ER methods only.
Some CIL methods use an online learning setting, where each training example is seen only once. In this paper, we primarily use offline methods. Note that while few-shot incremental learning [24] may appear similar to our work on class-incremental learning (CIL) with a small memory buffer, there is a significant difference. We assume that the data for new tasks is initially sufficient for effective learning, whereas few-shot incremental learning inherently deals with a scarcity of labeled samples.
Summary of contribution    We present TEAL, a novel strategy designed to select effective exemplars for small memory buffers. TEAL can be seamlessly integrated with replay-based class-incremental learning methods, significantly enhancing their performance.
2 Problem formulation
CIL involves learning from a sequence of disjoint tasks, where each task requires the classification of a set of classes, and every class belongs exclusively to a single task. At each incremental step, the learner receives the data of the new task and, in some cases, has access to a memory buffer containing exemplars from previous tasks. After each step, the learner is tested on the successful classification of all the classes it has seen, including those seen in previous steps.
More formally [see 19], a class-incremental learning problem $\mathcal{T}$ consists of a sequence of $T$ tasks. Each task $t$ contains labeled samples from $n_{t}$ different classes $X^{1},\dots,X^{n_{t}}$, where $X^{i}=\{(x_{1},i),\dots,(x_{m^{i}},i)\}$. The tasks do not share classes, i.e., $X^{i}\cap X^{j}=\emptyset~{}\forall i\neq j$. We denote by $N^{t}$ the total number of classes in all tasks up to task $t$: $N^{t}=\sum_{i=1}^{t}{n_{i}}$.
We consider the scenario where there is access to a memory buffer ${\mathcal{M}}$, which has a fixed size throughout the learning of $\mathcal{T}$. An incremental learner is a learning model (we consider only deep neural networks) that is trained sequentially on the tasks. Accordingly, the training on task $t$ is performed on data $X^{1},\dots,X^{n_{t}},{\mathcal{M}}$, where ${\mathcal{M}}$ contains exemplars from tasks $11,\dots,t-$ 1.
Different CIL methods may employ different loss functions, but at test time, they are all required to classify a given example into its predicted class while considering all previously seen classes.
3 Our method: TEAL
In active learning, it has been shown (empirically and formally) that when the number of labeled examples is small, it is best to choose the most typical examples for training. In the context of CIL and when the memory buffer is too small to truly represent the distribution of each class, we propose to adopt this strategy for the selection of points that are intended to populate the replay memory buffer. In other words, when the buffer is considerably small, it should contain representative exemplars and thus retain a more significant fraction of the previously acquired knowledge.
Our proposed method TEAL, Typicality Election Approach to continual Learning, aims to address only one component of the general Incremental Learning (IL) problem, namely, how to populate the memory buffer at the end of each IL iteration. Accordingly, after task $t$, TEAL should select for each seen class $n=\frac{|{\mathcal{M}}|}{N_{t}}$ exemplars to populate memory buffer ${\mathcal{M}}$.
Ideally, any new mechanism for memory population can be combined with any competitive Experience-Replay-Incremental-Learning (ER-IL) method in which this mechanism is a separate module. In accordance, we evaluate TEAL by considering alternative ER-IL methods, where we replace their native mechanism for buffer population by our proposed method. In almost all cases, this mechanism involves random sampling from the known labeled set, in order to faithfully represent the class distribution. In our evaluation study, we also consider an alternative population mechanism called Herding [22, 12].
3.1 Description of method
We begin with a definition of point typicality.
Definition 3.1 (Typicality).
For each exemplar $x$, let
$Typicality(x)=\left(\frac{1}{K}\sum_{x_{i}\in K\textrm{-}NN(x)}{\|x-x_{i}\|_{2}}\right)^{-1}$
where $K$ is a fixed number of nearest neighbors [^1] We use $K=20$, but other options yield similar results..
The essence of TEAL  is to populate the memory buffer with points that are both typical (as captured by the definition above) and diverse.
After training on task $t$, the incremental learner has to update the memory buffer ${\mathcal{M}}$ with data from the new classes $\{X^{N^{t-1}+1},\dots,X^{N^{t}}\}$. Given the fixed size of ${\mathcal{M}}$, the learner must first reduce the number of exemplars already in the buffer to accommodate new ones from the new classes. For each class, this involves repeatedly removing examples from the buffer after each IL step.
To manage this, TEAL  maintains a list of the selected exemplars from each class ordered by typicality, reflecting the order in which they are to be deleted from ${\mathcal{M}}$. This allows the learner to remove the least typical points in the buffer, whenever new classes appear. To achieve this goal, whenever a new class is introduced, TEAL selects the initial set of representative exemplars through an iterative process. The selected list is ordered such that the most typical exemplars, which are to be retained for as long as possible, are positioned first, while those that will be removed first are positioned last.
More specifically, in each repeat of this iterative selection process and for each class separately, TEAL selects a small number of exemplars (a fraction of $n$, where $n$ denotes the number of exemplars assigned to the respective class). Let $S_{i}$ denote the set of exemplars after the $i^{th}$ iteration, for a total of $k$ iterations. Note that $S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{k}$ and that $|S_{k}|=n$. The sizes $|S_{1}|,\dots,|S_{k}|$ define the selection pace, indicating the pace of selecting exemplars from new class data.
To initialize the selection process, we need a suitable embedding space for the class exemplars. To this end we train a deep model on all the available training data, and use activations in its penultimate layer as a representation for the new classes [cf. 22]. Subsequently, the selection process in the $i^{th}$ iteration involves 2 steps (see pseudocode in Alg. 1):
Step 1: Clustering
Since typical examples tend to cluster in bunches [10], and in order to ensure diversity, we seek typical exemplars from different regions of the embedding space. We achieve this by dividing the set of labeled points into $|S_{i}|$ clusters, where $|S_{i}|$ represents the number of exemplars that should be selected from this class after the $i^{th}$ iteration.
Step 2: Typicality
We then select the most typical point from each of the $|S_{i}|-|S_{i-1}|$ largest uncovered clusters, where an uncovered cluster is a cluster from which no point has already been selected. We order the selected points by their typicality score.
  Input: $X^{c}$eqparboxCOMMENT# a set of exemplars from class $c$
  Input: $n$eqparboxCOMMENT# number of examples to choose
  Require: $\phi:{\mathcal{X}}\to{\mathbb{R}}^{d}$eqparboxCOMMENT# current feature function
  Require: $s_{1},\dots,s_{k}$eqparboxCOMMENT# iterations pace, $s_{1}\leq\dots\leq s_{k}=n$
  Output: ${\mathcal{M}}_{c}$eqparboxCOMMENT# a set of n exemplars chosen from $X^{c}$
  $X^{c}_{emb}\leftarrow\phi(X^{c})$
  ${\mathcal{M}}_{c}\leftarrow\emptyset$
  for all $i=1,\dots,k$ do
     $C_{1},\dots,C_{s_{i}}\leftarrow$ clustering_algorithm$(X^{c}_{emb},\ s_{i})$eqparboxCOMMENT# $|C_{1}|\geq\dots\geq|C_{s_{i}}|$
     for all $j=1,\dots,s_{i}-s_{i-1}$ do
        add $\underset{x\in C_{j}}{\mathrm{argmax}}\,\{Typicality(x)\}$ to ${\mathcal{M}}_{c}$
     end for
  end for
  return ${\mathcal{M}}_{c}$
Algorithm 1 TEALConstructExemplarSet
3.2 Using TEAL for incremental learning
The integration of TEAL into a general class-incremental algorithm is described in Algs. 2-3.
  Input: ${\mathcal{M}}_{c}=(m_{1},\dots,m_{|{\mathcal{M}}_{c}|})$eqparboxCOMMENT# current exemplars set
  Input: $n$eqparboxCOMMENT# number of examples to keep
  return $(m_{1},\dots,m_{n})$
Algorithm 2 ReduceExemplarSet
  Input: $X^{t}=\{X^{N^{t-1}+1},\dots,X^{N^{t}}\}$eqparboxCOMMENT# training examples in per-class sets
  Require: $\theta^{t-1}$eqparboxCOMMENT# current model parameters
  Require: ${\mathcal{M}}=({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t-1}})$eqparboxCOMMENT# current exemplar sets
  $\theta^{t}\leftarrow\textrm{training\_model}(X^{t},{\mathcal{M}};\theta^{t-1})$
  $n\leftarrow\frac{|{\mathcal{M}}|}{N^{t}}$eqparboxCOMMENT# number of exemplars per class
  for $c\in\{1,\dots,N^{t-1}\}$ do
     ${\mathcal{M}}_{c}\leftarrow ReduceExemplarSet({\mathcal{M}}_{c},n)$
  end for
  for $c\in\{N^{t-1}+1,\dots,N^{t}\}$ do
     ${\mathcal{M}}_{c}\leftarrow ConstructExemplarSet(X^{c},n,\theta^{t})$
  end for
  ${\mathcal{M}}\leftarrow({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t}})$
Algorithm 3 IncrementalTrain
4 Empirical evaluation
We report two settings: (i) Integrated:we evaluate the beneficial contribution of TEAL to competitive ER-IL methods, by replacing their native selection strategy with TEAL (Section 4.2.1); (ii) Alternative selection strategies:we evaluate the beneficial contribution of TEAL as compared to Herding (Section 4.2.2). In both cases, we incrementally train a deep-learning model with a fixed-size replay buffer and monitor the average accuracy defined below upon completion of each task $t$.
4.1 Methodology
The majority of the experiments are conducted using the open-source CL library Avalanche [16], with the exception of those involving XDER [1], which is not integrated into Avalanche. For XDER, we utilize the code provided by the authors. In order to guarantee a fair comparison and in all experiments, we employ identical network architectures and maintain consistent experimental conditions.
We make certain in both settings that the buffer remains class-balanced, namely, each update maintains an equal representation of exemplars across all classes. When the buffer size is not evenly divisible by the number of classes, there may be instances where certain classes possess one additional exemplar. To maintain this balance, we follow this procedure: Prior to adding new exemplars we calculate $n=\frac{|{\mathcal{M}}|}{N^{t}}$, where $N^{t}$ is the total number of existing and new classes. Subsequently, we adjust the number of exemplars from existing classes to accommodate $n$ by removing the least likely points, followed by the addition of $n$ exemplars from each new class.
In the second setting, we employ a simple baseline ER-IL model. This model updates a fixed-size buffer of exemplars after training each task $t$ and replays it while training on task $t+1$. Additionally, we utilize a weighted data loader to ensure a balanced mix of data from both the new classes and the exemplars stored in the buffer in each batch. Across experiments, the only variation lies in the selection strategy employed: some experiments utilize random-sampling, others utilize Herding, and the remaining experiments leverage TEAL. In some experiments, we replace the native buffer selection strategy of an existing ER-IL method with TEAL or Herding to assess their impact. Other than this change in the method’s buffer population mechanism, everything else remains the same.
Datasets    We use several well known Continual Learning benchmarks. (i) Split CIFAR-100[22, 5], a dataset created by splitting CIFAR-100 [13] into 10 tasks, each containing 10 different classes of 32X32 images, with 500 images per class for training and 100 for testing. (ii) Split tinyImageNet, a dataset created by splitting tinyImageNet [14] into 10 tasks, each containing 20 different classes of 64X64 images, with 500 images per class for training and 50 for testing. (iii) Split CUB-200[3], a dataset created by splitting the CUB-200 [26] high-resolution image classification dataset, consisting of 200 categories of birds, with around 30 images per class for training and 30 for testing.
Architectures    In the first setting, except for one experiment, we employ a ResNet-18 model in all our experiments. The exception is the experiment involving the training of XDER on the Split CUB-200 dataset. Due to computational limitations, we used instead a pre-trained ResNet-50 model [11], and this was maintained under all the relevant conditions. In the second setting, we use a smaller version of ResNet-18 [11] as a simple baseline ER-IL model [see 17].
Metrics    As customary, we employ a score that reflects the accuracy during each stage of the incremental training. Let $T$ denote the total number of tasks, and $a_{t,i}$ denote the accuracy of task $i$ at the end of task $t$ ( $i\leq t\leq T$). For every task $t={1,\dots,T}$, define it average accuracy $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i}$. This metric provides a single value for each incremental step, enabling us to directly compare different methods at each step. Since all of the experiments we conduct are task-balanced in the sense that each task has an equal number of classes, there is no need to add a weighting term to the metric.
Baselines    The following methods are suitable for our investigation as they are designed to address ER-IL: XDER [1] updates the memory buffer by adding current information to past memories and preparing the model for new tasks. ER-ACE [2], or ER with Asymmetric Cross-Entropy, uses a separate loss for new and past tasks. BiC [28] introduces a bias correction layer to address class imbalance in incremental learning. iCaRL [22] combines the incremental classifier and representation learning with a nearest-mean-of-exemplars strategy, and Herding for exemplar selection. Since the classifier is based on the nearest-mean-of-exemplars, Herding is crucial for the selection strategy, as it guarantees that the features of the exemplars in the buffer approximate the original classes means. GEM [17] uses Gradient Episodic Memory to prevent gradient leakage between new and past tasks through constrained optimization. GDumb [21] employs a greedy sampling strategy for memory storage and retrains the neural network from scratch during inference. ER [4] explores different selection strategies in a simple ER framework.
In order to evaluate the competitiveness of these methods, we conducted multiple runs on Split CIFAR-100, utilizing fixed buffer sizes of 300, 500, and 2000. Fig. 2 shows the results.
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Figure 2: Baseline CIL methods comparison on Split CIFAR-100 with different fixed buffer sizes. For each method, each point represents the average accuracy achieved on the corresponding task $t$, $A_{t}$.
Probably because the memory buffer is so small, and as can be seen in Fig. 2, GEM and GDumb perform poorly in a relative comparison. BiC performs well on buffer size 2000, but significantly degrades for smaller buffers for which it is not suited. iCarl demonstrates high performance, but it is highly non-modular, being heavily dependent on the Herding selection strategy while using a K-Means classifier. Hence we cannot integrate it with TEAL.
We are therefore left with methods XDER, ER-ACE and ER, which are both suitable and competitive, to be examined with and without TEAL as the selection mechanism. For ER-ACE, since it depends on populating the buffer with exemplars at the beginning of the task before training on the new classes, we cannot rely on the representation provided by a trained model. Therefore, we adjust the integration of TEAL by filling the buffer in two stages: First, at the beginning of task $t$, we update the buffer with exemplars from the new classes using the random-sampling selection strategy. Second, after training task $t$, we replace the exemplars from the new classes in the buffer with exemplars chosen using TEAL. It’s important to note that throughout this process, the buffer maintains the fixed size of exemplars it allowed, ensuring compliance with the class-incremental framework.
4.2 Main results
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Split CIFAR-100
(a) $|{\mathcal{M}}|=300$
(e) $|{\mathcal{M}}|=1000$
(f) $|{\mathcal{M}}|=2000$
Split tinyImageNet
(d) $|{\mathcal{M}}|=400$
(h) ER
(i) XDER
Split CUB-200
(g) ER-ACE
Figure 3: The performance of ER-IL methods with and without TEAL. Each plot displays the average accuracy after training incrementally on a different number of classes. Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy. The error bars correspond to standard error based on 4-10 repetitions.
4.2.1 TEAL integrated with SOTA ER-IL methods
Split CIFAR-100
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
(e) $|{\mathcal{M}}|=1000$
(f) $|{\mathcal{M}}|=2000$
Split tinyImageNet
(d) $|{\mathcal{M}}|=400$
(h) ER
(i) XDER
Split CUB-200
(g) ER-ACE
Figure 4: Performance improvement of TEAL when integrated with different ER-IL methods. Each graph depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks. The first two rows present the three methods improvement on Split CIFAR-100 and Split tinyImageNet with different fixed buffer sizes denoted as $|{\mathcal{M}}|$. The last row pertains to Split CUB-200, where each graph showcases different buffer sizes for a different method.
Table 1: Final average accuracy for method with and without TEAL
Dataset Split CIFAR-100 Split tinyImageNet Split CUB-200
$|{\mathcal{M}}|$ 300 500 2000 400 1000 2000 200 400
XDER 41.97±0.25 47.97±0.22 57.69±0.21 20.49±0.19 32.68±0.24 39.76±0.31 43.63±1.78
XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17 23.61±0.32 34.28±0.24 40.76±0.23 49.55±0.47
Improvement 3.08 2.32 1.1 3.13 1.59 1.0 5.96
ER-ACE 35.63±0.36 40.6±0.34 53.6±0.17 13.02±0.2 17.37±0.34 21.71(±0.17) 8.56±0.15 11.25±0.99
ER-ACE+TEAL 40.59±0.2 44.25±0.35 56.11±0.19 14.43±0.29 18.88±0.21 24.12±0.24 10.41±0.16 12.33±0.26
Improvement 4.97 3.65 2.51 1.41 1.51 2.4 1.85 1.08
ER 14.82±0.21 18.21±0.29 36.54±0.31 7.75±0.06 7.7±0.06 8.27±0.08 8.98±0.18 12.01±0.52
ER+TEAL 17.97±0.25 22.39±0.31 38.17±0.37 7.86±0.05 7.99±0.06 8.82±0.09 10.37±0.23 14.13.76±0.3
Improvement 3.15 4.18 1.62 0.11 0.29 0.56 1.39 2.12
As mentioned above, we investigate the baseline methods XDER, ER-ACE and ER, comparing their performance when using either their native selection strategies or TEAL. We examined Split CIFAR-100 with buffer sizes of 300, 500 and 2000, Split tinyImageNet with buffer sizes of 400, 1000 and 2000, and Split CUB-200 with buffer sizes of 200222Due to computational limitation of the relevant package, we did not run XDER with buffer size 200. and 400. Fig. 3 shows the average accuracy $A_{t}$ after each task $t$. Fig. 4 shows the corresponding improvement of TEAL. Finally, the results of the final average accuracy $A_{t}$ and the differences ’improved-vanilla’ are reported in Table 1.
As demonstrated in Fig. 4, in most cases, the improvement achieved by TEAL increases as incremental training progresses, resulting in a more significant improvement by the final task. Indeed, we do not expect significant differences immediately after the first task, as catastrophic forgetting hasn’t occurred yet. Reassuringly, for the majority of the experiments, the improvement becomes apparent starting from the second task.
Table 1 shows that the difference (’improvement’ row) is always positive, namely, TEAL always improves the outcome. Looking at the improvement row for the XDER method, it is evident that the improvement provided by TEAL increases as the buffer size decreases. This trend is also demonstrated for ER-ACE and partially for ER on Split CIFAR-100.
4.2.2 Comparison of selection strategy
We begin with a simple ER-IL model on Split CIFAR-100333The order of the classes and the partition to tasks is randomly selected and fixed for all comparisons. with various buffer sizes, and three different selection strategies: random, Herding and TEAL. As depicted in Fig. 5, our method enhances performance as compared to random selection by almost 3% for smaller buffers (300, 400, 500) and about 2.5% for larger buffers (1000, 2000). When compared to Herding, our method is more effective for smaller buffer sizes, enhancing performance by up to 1.2% for a buffer size of 300, but less effective with large buffers. Similar results, using different class ordering, are shown in App A.1.
Figure 5: Split CIFAR-100: performance difference between TEAL and either random-sampling or Herding selection. Each bar corresponds to a fixed buffer size.
Dataset Split CIFAR-100
$|{\mathcal{M}}|$ 300 500 2000
XDER+TEAL_OneTime 44.13±0.26 49.87±0.22 59.05±0.14
XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17
ER-ACE+TEAL_OneTime 35.81±0.42 40.16±0.3 51.52±0.24
ER-ACE+TEAL 36.4±0.12 40.26±0.23 51.68±0.35
ER+TEAL_OneTime 17.64±0.12 22.01±0.18 40.65±0.32
ER+TEAL 18.1±0.1 22.07±0.13 39.96±0.23
Table 2: The improvement of 2 variants of TEAL when integrated with different CIL methods: XDER, ER-ACE and ER. In almost all cases, the original TEAL matches or outperforms the OneTime variant.
Given the results shown in Fig. 5, wherein TEAL seems less effective than herding with large buffers when training the simple baseline model, we repeat this comparison while integrating both methods - TEAL and Herding - into competitive ER-IL methods. Fig. 6 shows the difference between the improvement obtained by TEAL and the one obtained by Herding. Since in almost all cases this difference is positive, we may conclude that TEAL is significantly more beneficial than Herding.
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Figure 6: The difference between the improvement obtained by TEAL and the one obtained by Herding, showing 2 ER-IL methods and 3 buffer sizes while training on the Split CIFAR-100 dataset.
4.3 Ablation study
We explore an alternative variant of TEAL, which selects the set of exemplars in a single pass. We call this variant TEAL_OneTime. While still selecting exemplars for each class separately, this variant first partitions the exemplars of each class into $n=\frac{|{\mathcal{M}}|}{N^{t}}$ clusters, and then selects the most typical exemplar from each cluster in descending order of cluster size, preserving the order of selections.
The experiments on the two variants are conducted using Split CIFAR-100 with the smaller version of ResNet-18 mentioned in Section 4.1 and buffer sizes of 300, 500 and 2000. We investigate the performance enhancement of each variant when integrated with XDER, ER-ACE, and ER. The results are shown in Table 2. Whenever there is a significant difference in performance, the iterative variant outperforms the other one in 4 out of 5 cases. However, the differences are not substantial.
5 Summary and discussion
We propose a new mechanism to select exemplars for the memory buffer in replay-based CIL methods. This method is based on the principles of diversity and representativeness. When the memory buffer is relatively small, our method TEAL is shown to outperform both the native mechanism of each ER-IL method (usually random selection) and an alternative selection mechanism called Herding. Even when the buffer is large, our method is beneficial in almost all studied cases.
Our future work will investigate ways to determine at which point, if any, the added value of our method diminishes or possibly becomes harmful. We will also integrate TEAL with other replay-based CIL methods that utilize different selection strategies, and analyze the impact on their performance.
Acknowledgement
This work was supported by grants from the Israeli Council of Higher Education and the Gatsby Charitable Foundations.
Appendix
Appendix A Implementation details
The source code used in this study would be uploaded to our GitHub repository and will be made public upon acceptance.
TEAL, as described in Alg. 1, requires an iterations pace $s_{1}\leq\dots\leq s_{k}=n$ which indicates the pace of selecting exemplars from new class data. In both settings we use a logarithmic pace. We set a base $b=1.4$, and define $s_{1}=\lfloor b^{4}\rfloor,s_{2}=\lfloor b^{5}\rfloor,\dots,s_{k}=\lfloor b^{log_{b}n}\rfloor$.
A.1 Stand-alone setting
For all selection strategies, we use a smaller ResNet-18, as mentioned above, trained for 200 epochs. Optimization is performed with an SGD optimizer using Nesterov momentum of 0.9, a weight decay of 0.0002, and a learning rate that starts at 0.1 and decays by a factor of 0.3 every 66 epochs. Training is conducted with a batch size of 128 examples, and data augmentation is applied through random cropping and horizontal flips.
We run this setting on Split CIFAR-100 with three different random orders of classes:
  1. 1.
    The Order we display on Fig. 5 is: [44, 19, 93, 90, 71, 69, 37, 95, 53, 91, 81, 42, 80, 85, 74, 56, 76, 63, 82, 40, 26, 92, 57, 10, 16, 66, 89, 41, 97, 8, 31, 24, 35, 30, 65, 7, 98, 23, 20, 29, 78, 61, 94, 15, 4, 52, 59, 5, 54, 46, 3, 28, 2, 70, 6, 60, 49, 68, 55, 72, 79, 77, 45, 1, 32, 34, 11, 0, 22, 12, 87, 50, 25, 47, 36, 96, 9, 83, 62, 84, 18, 17, 75, 67, 13, 48, 39, 21, 64, 88, 38, 27, 14, 73, 33, 58, 86, 43, 99, 51]
  2. 2.
    The order in 7(a) is: [45, 15, 90, 32, 35, 63, 17, 72, 79, 96, 48, 36, 16, 11, 23, 80, 22, 58, 3, 62, 50, 33, 66, 99, 43, 76, 7, 57, 81, 82, 6, 10, 24, 52, 95, 73, 91, 21, 38, 31, 85, 59, 13, 69, 75, 70, 64, 8, 77, 34, 46, 39, 92, 0, 44, 98, 49, 9, 4, 61, 12, 83, 28, 78, 40, 88, 54, 5, 26, 41, 89, 20, 84, 2, 1, 55, 19, 74, 25, 37, 42, 14, 30, 18, 67, 71, 68, 27, 60, 51, 29, 56, 93, 47, 97, 94, 86, 87, 65, 53]
  3. 3.
    The order in 7(b) is: [48, 97, 1, 81, 90, 49, 10, 8, 7, 20, 70, 73, 75, 14, 91, 38, 47, 21, 74, 52, 80, 98, 59, 12, 71, 85, 6, 34, 55, 82, 95, 63, 78, 15, 94, 60, 99, 76, 25, 40, 88, 0, 62, 96, 87, 51, 16, 18, 9, 19, 29, 45, 86, 53, 56, 31, 28, 61, 30, 33, 4, 67, 64, 58, 50, 54, 3, 13, 37, 27, 66, 77, 84, 69, 2, 41, 22, 92, 42, 44, 11, 36, 46, 79, 65, 72, 23, 17, 39, 5, 89, 35, 24, 83, 43, 57, 93, 32, 68, 26]
(a) Order 2
(b) Order 3
Figure 7: Enhanced accuracy of TEAL compared to random-sampling and Herding selection strategies, additional two different classes order.
A.2 Integrated setting
Here we train a ResNet-18 model for 100 epochs using the same optimizer as in A.1, with the same batch size and data augmentations, with some exceptions. ER-ACE starts with a learning rate of 0.01, and all experiments on Split CUB-200 are conducted for 30 epochs with a batch size of 16 due to the dataset size and the resolution of its images.
A.3 Baselines setting
The setting for the experiments of the baseline methods is the same as in A.2 with some exceptions. For BiC, the number of training epochs is 250, and the learning rate scheduler decays the learning rate by a factor of 0.1 on epochs 100, 150 and 200. For GEM, the batch size is 32 and the learning rate starts from 0.03.For GDumb, the batch size is 32. For iCaRL the learning rate starts from 2, the weight decay is 0.00001 and the learning rate scheduler decays the learning rate by a factor of 0.2 on epochs 49 and 63.
Appendix B Compute resources
All experiments involved training deep learning models, necessitating the use of GPUs. For the Split CIFAR-100 experiments, 10 GB of GPU memory was used. The Split tinyImageNet experiments required 22 GB of GPU memory, while the Split CUB-200 experiments utilized 45 GB of GPU memory. Any other experiments conducted for the full research and not reported in the paper required the same compute resources.

### v2

@@
-TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremenal Learning
+TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremental Learning
 Abstract
 Continual Learning is an unresolved challenge, whose relevance increases when considering modern applications.
-Unlike the human brain, trained deep neural networks suffer from a phenomenon called Catastrophic Forgetting, where they progressively lose previously acquired knowledge upon learning new tasks.
-To mitigate this problem, numerous methods have been developed, many relying on replaying past exemplars during new task training.
+Unlike the human brain, trained deep neural networks suffer from a phenomenon called catastrophic forgetting, wherein they progressively lose previously acquired knowledge upon learning new tasks.
+To mitigate this problem, numerous methods have been developed, many relying on the replay of past exemplars during new task training.
 However, as the memory allocated for replay decreases, the effectiveness of these approaches diminishes.
 On the other hand, maintaining a large memory for the purpose of replay is inefficient and often impractical.
-Here we introduce TEAL, a novel approach to populate the memory with exemplars, that can be integrated with various experience-replay methods and significantly enhance their performance on small memory buffers.
-We show that TEAL improves the average accuracy of the SOTA method XDER as well as ER and ER-ACE on several image recognition benchmarks, with a small memory buffer of 1-3 exemplars per class in the final task.
-This confirms the hypothesis that when memory is scarce, it is best to prioritize the most typical data.
+Here we introduce TEAL, a novel approach to populate the memory with exemplars, that can be integrated with various experience-replay methods and significantly enhance their performance with small memory buffers.
+We show that TEAL enhances the average accuracy of existing class-incremental methods and outperforms other selection strategies, achieving state-of-the-art performance even with small memory buffers of 1-3 exemplars per class in the final task.
+This confirms our initial hypothesis that when memory is scarce, it is best to prioritize the most typical data.
+Code is available at this https URL.
 1 Introduction
-With the recent advances in deep neural networks, research interest in incremental learning has increased.
-Here, the problem of catastrophic forgetting, as defined by McCloskey and Cohen [20], can be severe.
-The necessity to incorporate new task knowledge into an already trained network has become more relevant given that training over extensive datasets is time-consuming.
-It is often impractical to retrain the network from scratch on the combined original and new task data.
-Furthermore, access to the already-trained data might be limited or non-existent.
-Figure 1: Catastrophic forgetting when training CIFAR-10 [13] sequentially with only finetuning.
-Various methods have been developed to tackle this challenge across different frameworks, categorized by Van de Ven et al.
-[25] into three types: task-incremental, domain-incremental, and class-incremental learning.
-The core concept of incremental learning is that a model must learn tasks sequentially, one after another.
-Among these, class-incremental learning is considered the most challenging.
-In this scenario, each task introduces new classes, and the model must identify the specific class of each input without access to the task ID on which it was trained.
-There are several approaches to incremental learning, each with its own distinct assumptions and configurations.
-In this paper, we adhere to the constrained setup described by De Lange et al.
-[7], which does not rely on task boundaries during training or testing.
-This setup maintains a constant memory size throughout incremental training, ensuring that it does not exceed a predetermined limit established at the outset.
-At each incremental step, we can add new exemplars to the memory buffer only after space has been allocated for them.
+With the recent advances in deep neural networks, there has been a growing research interest in incremental learning.
+The need to integrate new task knowledge into an already trained network has become increasingly important, especially considering the time-consuming nature of training on large datasets.
+Retraining the network from scratch on both the original and new task data is often impractical, and access to the original training data may be limited or unavailable.
+In this context, catastrophic forgetting, as described by McCloskey and Cohen (1989), can be particularly severe.
+To address this challenge, various methods have been developed across different frameworks.
+Van de Ven, Tuytelaars, and Tolias (2022) categorizes these methods into three types: task-incremental, domain-incremental, and class-incremental learning.
+The fundamental idea behind incremental learning is that a model must sequentially learn tasks, one after the other.
+Among these approaches, Class-Incremental Learning (CIL) is recognized as the most challenging.
+Here, each task introduces new classes, and the model must accurately identify the class of each input without access to the corresponding task ID, see Fig.
+1.
+Figure 1: Illustration of CIL with Experience Replay.
+Incremental learning can be approached in various ways, each with its own assumptions and configurations.
+In this paper, we follow the constrained setup outlined by De Lange et al.
+(2021), which does not depend on task boundaries during either training or testing.
+This approach maintains a fixed memory size throughout incremental training, ensuring it stays within a predefined limit set from the beginning.
+At each incremental step, new exemplars can be added to the memory buffer only after sufficient space has been vacated.
 We focus our attention on a prevalent and rather successful framework called Experience Replay (ER), which involves storing a set of exemplars in memory and reusing them for rehearsal purposes while training on new tasks.
 Within this framework, several strategies exist for selecting which exemplars to retain in memory.
-Masana et al.
-[19] demonstrate that the most successful strategies are either random-sampling or Herding as defined by Welling [27].
 Not surprisingly, the smaller the memory buffer is, the less effective the strategy is at mitigating catastrophic forgetting.
 This leads us to the following question: Are these strategies necessarily optimal for all memory sizes?
-In other words, could different strategies prove to be more effective depending on the size of the memory buffer?
-Our proposed method, TEAL, operates on similar principles, aiming to identify a set of representative exemplars that also exhibit diversity.
-An exemplar is deemed representative if its likelihood, when considered within the distribution of all points in an appropriate latent space, is high.
-Leveraging clustering ensures the selection of a diverse set of exemplars.
-This approach enhances the model’s capacity to retain previous knowledge while effectively learning new tasks, even when constrained by a small memory buffer.
-Central to the success of this approach is our ability to derive an appropriate latent space.
-The subsequent sections elaborate on our methodology, experimental setup, comparative performance against alternative selection strategies, and the enhancements TEAL offers to various existing class-incremental methods.
-Related work    Several approaches exist for incremental learning [see 7].
-Alongside ER, alternatives include Generative Replay (GR) [23], Parameter Isolation, and Regularization.
-Similar to ER, GR replays data from previous tasks during new task training, but uses a generative model to create new samples instead of retaining exemplars seen by the model [6, 8, 9].
-Parameter Isolation assigns distinct parameters to each task to reduce forgetting [18] by fixing parameters assigned to previous tasks, while Regularization-based methods [15] incorporate additional terms into the loss function to retain prior knowledge while learning from new data.
-These methods differ in how they utilize memory: ER stores exemplars, GR stores a generative model, and Parameter Isolation stores task-specific parameters.
+In other words, can different strategies be found suitable for different sizes of the memory buffer?
+In active learning, it has been shown (empirically and formally) that when the number of labeled examples is small, it is best to choose the most typical examples for training (Hacohen, Dekel, and Weinshall 2022).
+In the context of CIL and when the memory buffer is too small to truly represent the distribution of each class, we propose to adopt this strategy for the selection of points that are intended to populate the replay memory buffer.
+In other words, when the buffer is considerably small, it should contain representative exemplars and thus retain a more significant fraction of the previously acquired knowledge.
+Our proposed method TEAL, Typicality Election Approach to continual Learning, is primarily targeted at scenarios with small buffers.
+While the use of small buffers may seem too strict, there are situations where maintaining a large memory for replay is simply not feasible.
+For example, applications running on mobile devices face severe memory constraints and could greatly benefit from methods specifically designed for small buffers.
+Accordingly, TEAL  aims to identify a set of representative exemplars that also exhibit diversity.
+An exemplar is deemed representative if its likelihood, when considered within the distribution of all points, is high.
+To ensure diversity, data clustering is leveraged.
+Central to the success of our approach is the ability to derive an appropriate latent space, where data clustering and likelihood estimation can be reliably obtained.
+Ideally, any mechanism for memory population can be combined with any competitive Experience-Replay-Incremental-Learning (ER-IL) method in which this mechanism is a separate module.
+In accordance, we evaluate our approach in Section 4 by considering alternative ER-IL methods, where we replace their native mechanism for buffer population by TEAL.
+The method, methodology and its experimental evaluation are described in the rest of this paper, with emphasis on the enhancements TEAL offers to various existing class-incremental methods.
+Related work    Several approaches exist for incremental learning (see De Lange et al.
+2021), including Experience Replay (ER), Generative Replay (GR) (Shin et al.
+2017), Parameter Isolation, and Regularization.
+Similar to ER, GR replays data from previous tasks during new task training, but uses a generative model to create new samples instead of retaining exemplars seen by the model (Choi, El-Khamy, and Lee 2021; Gao and Liu 2023; Gautam et al.
+2024).
+Parameter Isolation assigns distinct parameters to each task to reduce forgetting (Mallya and Lazebnik 2018) by fixing parameters assigned to previous tasks, while Regularization-based methods (Li and Hoiem 2017) incorporate additional terms into the loss function to retain prior knowledge while learning from new data.
+These methods differ in the ways they utilize memory: ER stores exemplars, GR stores a generative model, and Parameter Isolation stores task-specific parameters.
 Regularization-based methods do not rely on memory.
 This diversity makes it challenging to compare methods directly.
-In particular, we note that since generative models tend to be very large, this approach is hardly suitable to the domain of small memory buffers addressed below.
-Hence, our focus in this paper is on ER methods only.
-Some CIL methods use an online learning setting, where each training example is seen only once.
-In this paper, we primarily use offline methods.
-Note that while few-shot incremental learning [24] may appear similar to our work on class-incremental learning (CIL) with a small memory buffer, there is a significant difference.
-We assume that the data for new tasks is initially sufficient for effective learning, whereas few-shot incremental learning inherently deals with a scarcity of labeled samples.
+In particular, we note that since generative models tend to be very large, GR is hardly suitable to the domain of small memory buffers addressed here.
+A similar concern can be raised concerning Parameter Isolation methods.
+Hence, our focus in this paper is on ER.
+Note that while few-shot incremental learning (Tian et al.
+2024) may appear similar to our work on CIL with a small memory buffer, there is a significant difference.
+Specifically, we assume that the data for new tasks is initially sufficient for effective learning, whereas few-shot incremental learning inherently deals with a scarcity of labeled samples.
+When comparing strategies for the population of the memory buffer, Masana et al.
+(2022) demonstrate that the most successful strategies are either random-sampling or Herding as defined by Welling (2009).
+The latter strategy involves retaining a set of exemplars whose mean is closest to the class mean.
+Another successful strategy presented by Bang et al.
+(2021) leverages classification uncertainty and data augmentation to enhance the diversity of data instances (Uncertainty).
+Other selection strategies, such as GSS (Aljundi et al.
+2019) and selecting the exemplars with the highest entropy of the softmax outputs (Chaudhry et al.
+2018a), have been shown in previous studies to be inferior to random sampling and Herding (Masana et al.
+2022; Prabhu et al.
+2023).
+Therefore, we do not include them in our comparative empirical evaluation.
+Recently, Hacohen and Tuytelaars (2024) proposed a selection strategy called Goldilocks, which retains exemplars learned at an intermediate pace, accommodating various buffer sizes.
+Since the results presented in this paper are primarily within a task-incremental framework (which yields higher average accuracy) and the paper does not provide code, we do not include this method in our comparisons.
 Summary of contribution    We present TEAL, a novel strategy designed to select effective exemplars for small memory buffers.
-TEAL can be seamlessly integrated with replay-based class-incremental learning methods, significantly enhancing their performance.
+TEAL is seamlessly integrated with a number of SOTA replay-based class-incremental learning method, and significantly enhances their performance.
 2 Problem formulation
-CIL involves learning from a sequence of disjoint tasks, where each task requires the classification of a set of classes, and every class belongs exclusively to a single task.
-At each incremental step, the learner receives the data of the new task and, in some cases, has access to a memory buffer containing exemplars from previous tasks.
-After each step, the learner is tested on the successful classification of all the classes it has seen, including those seen in previous steps.
-More formally [see 19], a class-incremental learning problem $\mathcal{T}$ consists of a sequence of $T$ tasks.
-Each task $t$ contains labeled samples from $n_{t}$ different classes $X^{1},\dots,X^{n_{t}}$, where $X^{i}=\{(x_{1},i),\dots,(x_{m^{i}},i)\}$.
-The tasks do not share classes, i.e., $X^{i}\cap X^{j}=\emptyset~{}\forall i\neq j$.
+A CIL problem $\mathcal{T}$ consists of a sequence of $T$ tasks.
+Each task $t\in T$ contains a set of classes $C^{t}=(c^{1},\dots,c^{n_{t}})$ and labeled samples from these classes $X_{t}=(X^{1},\dots,X^{n_{t}})$, where $X^{i}=\{(x_{1},i),\dots,(x_{m^{i}},i)\}$.
+The tasks do not share classes, i.e., $C^{t_{i}}\cap C^{t_{j}}=\emptyset~{}\forall i\neq j$.
 We denote by $N^{t}$ the total number of classes in all tasks up to task $t$: $N^{t}=\sum_{i=1}^{t}{n_{i}}$.
@@
 An incremental learner is a learning model (we consider only deep neural networks) that is trained sequentially on the tasks.
-Accordingly, the training on task $t$ is performed on data $X^{1},\dots,X^{n_{t}},{\mathcal{M}}$, where ${\mathcal{M}}$ contains exemplars from tasks $11,\dots,t-$ 1.
-Different CIL methods may employ different loss functions, but at test time, they are all required to classify a given example into its predicted class while considering all previously seen classes.
+Accordingly, the training on task $t$ is performed on data $X_{t}\cup{\mathcal{M}}$, where ${\mathcal{M}}$ contains stored exemplars from tasks $11,\dots,t-$ 1.
+At test time, a CIL method is required to classify a given example into its predicted class while considering all previously seen classes.
+Our proposed method TEAL is described in Section 3.
+It aims to address only one component of the general Incremental Learning (IL) problem, namely, how to populate the memory buffer at the end of each IL iteration.
+Accordingly, after task $t$, TEAL should select for each seen class $n=\frac{|{\mathcal{M}}|}{N_{t}}$ exemplars to populate memory buffer ${\mathcal{M}}$.
 3 Our method: TEAL
-In active learning, it has been shown (empirically and formally) that when the number of labeled examples is small, it is best to choose the most typical examples for training.
-In the context of CIL and when the memory buffer is too small to truly represent the distribution of each class, we propose to adopt this strategy for the selection of points that are intended to populate the replay memory buffer.
-In other words, when the buffer is considerably small, it should contain representative exemplars and thus retain a more significant fraction of the previously acquired knowledge.
-Our proposed method TEAL, Typicality Election Approach to continual Learning, aims to address only one component of the general Incremental Learning (IL) problem, namely, how to populate the memory buffer at the end of each IL iteration.
-Accordingly, after task $t$, TEAL should select for each seen class $n=\frac{|{\mathcal{M}}|}{N_{t}}$ exemplars to populate memory buffer ${\mathcal{M}}$.
-Ideally, any new mechanism for memory population can be combined with any competitive Experience-Replay-Incremental-Learning (ER-IL) method in which this mechanism is a separate module.
-In accordance, we evaluate TEAL by considering alternative ER-IL methods, where we replace their native mechanism for buffer population by our proposed method.
-In almost all cases, this mechanism involves random sampling from the known labeled set, in order to faithfully represent the class distribution.
-In our evaluation study, we also consider an alternative population mechanism called Herding [22, 12].
-3.1 Description of method
+Figure 2: Illustration of TEAL’s iterative class selection process, which establishes a priority order for the selected set.
+Initially, an embedding space is generated separately for each class (shown on the left).
+Samples are then selected iteratively with $s_{1}=4$, $s_{2}=7$, and $s_{3}=9$ (see text for details).
+Each row on the right panel represents one iteration: the left image displays the $s_{i}$ clusters (obtained using K-means) with the previously selected points $S_{i-1}$ marked with ’o’, while the right image shows the updated set $S_{i}$, with newly selected samples marked with ’X’.
+In the last iteration 3 clusters remain uncovered, but only 2 samples are selected, leaving the red cluster uncovered by $S_{3}$.
 We begin with a definition of point typicality.
@@
 For each class, this involves repeatedly removing examples from the buffer after each IL step.
-To manage this, TEAL  maintains a list of the selected exemplars from each class ordered by typicality, reflecting the order in which they are to be deleted from ${\mathcal{M}}$.
-This allows the learner to remove the least typical points in the buffer, whenever new classes appear.
-To achieve this goal, whenever a new class is introduced, TEAL selects the initial set of representative exemplars through an iterative process.
-The selected list is ordered such that the most typical exemplars, which are to be retained for as long as possible, are positioned first, while those that will be removed first are positioned last.
-More specifically, in each repeat of this iterative selection process and for each class separately, TEAL selects a small number of exemplars (a fraction of $n$, where $n$ denotes the number of exemplars assigned to the respective class).
-Let $S_{i}$ denote the set of exemplars after the $i^{th}$ iteration, for a total of $k$ iterations.
-Note that $S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{k}$ and that $|S_{k}|=n$.
-The sizes $|S_{1}|,\dots,|S_{k}|$ define the selection pace, indicating the pace of selecting exemplars from new class data.
+To address this challenge, TEAL maintains a priority list of selected exemplars from each class, which reflects the order in which they should be removed from ${\mathcal{M}}$.
+This approach allows the learner to remove the least typical points from the buffer when new classes emerge.
+The list is structured so that the most typical and diverse exemplars, which should be retained as long as possible, are positioned at the top, while those slated for earlier removal are positioned at the bottom.
+More specifically, consider a fixed class, and let $n$ denote the number of exemplars assigned to that class.
+TEAL repeatedly selects a small fraction of $n$, generating a sequence of subsets $S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{k}$, where $s_{i}=|S_{i}|$ and $s_{1}\leq\dots\leq s_{k}=n$.
+The sizes ${s_{i}}$ define the selection pace, indicating the rate at which exemplars are accumulated.
+The inclusion relation induces the priority order: points in $S_{k}\setminus S_{k-1}$ are removed first, while points in $S_{1}$ are removed last.
+This process is illustrated in Fig.
+2.
 To initialize the selection process, we need a suitable embedding space for the class exemplars.
-To this end we train a deep model on all the available training data, and use activations in its penultimate layer as a representation for the new classes [cf.
-22].
+To this end we train a deep model on all the available training data, and use activations in its penultimate layer as a representation for the new classes.
 Subsequently, the selection process in the $i^{th}$ iteration involves 2 steps (see pseudocode in Alg.
 1):
-Step 1: Clustering
-Since typical examples tend to cluster in bunches [10], and in order to ensure diversity, we seek typical exemplars from different regions of the embedding space.
-We achieve this by dividing the set of labeled points into $|S_{i}|$ clusters, where $|S_{i}|$ represents the number of exemplars that should be selected from this class after the $i^{th}$ iteration.
-Step 2: Typicality
-We then select the most typical point from each of the $|S_{i}|-|S_{i-1}|$ largest uncovered clusters, where an uncovered cluster is a cluster from which no point has already been selected.
-We order the selected points by their typicality score.
-  Input: $X^{c}$eqparboxCOMMENT# a set of exemplars from class $c$
-  Input: $n$eqparboxCOMMENT# number of examples to choose
-  Require: $\phi:{\mathcal{X}}\to{\mathbb{R}}^{d}$eqparboxCOMMENT# current feature function
-  Require: $s_{1},\dots,s_{k}$eqparboxCOMMENT# iterations pace, $s_{1}\leq\dots\leq s_{k}=n$
-  Output: ${\mathcal{M}}_{c}$eqparboxCOMMENT# a set of n exemplars chosen from $X^{c}$
-  $X^{c}_{emb}\leftarrow\phi(X^{c})$
-  ${\mathcal{M}}_{c}\leftarrow\emptyset$
-  for all $i=1,\dots,k$ do
-     $C_{1},\dots,C_{s_{i}}\leftarrow$ clustering_algorithm$(X^{c}_{emb},\ s_{i})$eqparboxCOMMENT# $|C_{1}|\geq\dots\geq|C_{s_{i}}|$
-     for all $j=1,\dots,s_{i}-s_{i-1}$ do
-        add $\underset{x\in C_{j}}{\mathrm{argmax}}\,\{Typicality(x)\}$ to ${\mathcal{M}}_{c}$
-     end for
-  end for
-  return ${\mathcal{M}}_{c}$
-Algorithm 1 TEALConstructExemplarSet
-3.2 Using TEAL for incremental learning
-The integration of TEAL into a general class-incremental algorithm is described in Algs. 2-3.
-  Input: ${\mathcal{M}}_{c}=(m_{1},\dots,m_{|{\mathcal{M}}_{c}|})$eqparboxCOMMENT# current exemplars set
-  Input: $n$eqparboxCOMMENT# number of examples to keep
-  return $(m_{1},\dots,m_{n})$
-Algorithm 2 ReduceExemplarSet
-  Input: $X^{t}=\{X^{N^{t-1}+1},\dots,X^{N^{t}}\}$eqparboxCOMMENT# training examples in per-class sets
-  Require: $\theta^{t-1}$eqparboxCOMMENT# current model parameters
-  Require: ${\mathcal{M}}=({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t-1}})$eqparboxCOMMENT# current exemplar sets
-  $\theta^{t}\leftarrow\textrm{training\_model}(X^{t},{\mathcal{M}};\theta^{t-1})$
-  $n\leftarrow\frac{|{\mathcal{M}}|}{N^{t}}$eqparboxCOMMENT# number of exemplars per class
-  for $c\in\{1,\dots,N^{t-1}\}$ do
-     ${\mathcal{M}}_{c}\leftarrow ReduceExemplarSet({\mathcal{M}}_{c},n)$
-  end for
-  for $c\in\{N^{t-1}+1,\dots,N^{t}\}$ do
-     ${\mathcal{M}}_{c}\leftarrow ConstructExemplarSet(X^{c},n,\theta^{t})$
-  end for
-  ${\mathcal{M}}\leftarrow({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t}})$
-Algorithm 3 IncrementalTrain
+Step 1: Clustering.
+In order to ensure diversity, we seek typical exemplars from different regions of the embedding space.
+When constructing set $S_{i}$, we achieve this by dividing the set of labeled points into $s_{i}$ clusters using K-Means (Lloyd 1982).
+Step 2: Typicality.
+We then select the most typical point from each of the $s_{i}-s_{i-1}$ largest uncovered clusters, where an uncovered cluster is a cluster from which no point has already been selected.
+Algorithm 1 ConstructExemplarSet
+Input: a set of exemplars from class $c$ $X^{c}$, number of examples to choose $n$
+Require: current feature function $\phi:{\mathcal{X}}\to{\mathbb{R}}^{d}$, iterations pace $s_{1},\dots,s_{k}=n$
+Output: a set of n exemplars ${\mathcal{M}}_{c}\subseteq X^{c}$
+1:  $X^{c}_{emb}\leftarrow\phi(X^{c})$
+2:  ${\mathcal{M}}_{c}\leftarrow\emptyset$
+3:  for all $i=1,\dots,k$ do
+4:     $C_{1},\dots,C_{s_{i}}\leftarrow$ clustering_algorithm$(X^{c}_{emb},\ s_{i})$triangleright $|C_{1}|\geq\dots\geq|C_{s_{i}}|$
+5:     for all $j=1,\dots,s_{i}-s_{i-1}$ do
+6:        if $C_{j}$ is uncovered  then
+7:           add $\operatorname*{argmax}_{x\in C_{j}}\{Typicality(x)\}$ to ${\mathcal{M}}_{c}$
+8:        end if
+9:     end for
+10:  end for
+11:  return ${\mathcal{M}}_{c}$
+The integration of TEAL into a general class-incremental algorithm is described in Alg. 2.
+Algorithm 2 IncrementalTraining
+Input: training examples $X_{t}$
+Require: current exemplar sets ${\mathcal{M}}=({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t-1}})$, current model parameters $\theta^{t-1}$
+1:  $\theta^{t}\leftarrow\textrm{training\_model}(X_{t},{\mathcal{M}};\theta^{t-1})$
+2:  $n\leftarrow\frac{|{\mathcal{M}}|}{N^{t}}$
+3:  for $c\in\{1,\dots,N^{t-1}\}$ do
+4:     ${\mathcal{M}}_{c}\leftarrow$ first $n$ exemplars in ${\mathcal{M}}_{c}$
+5:  end for
+6:  for $c\in\{N^{t-1}+1,\dots,N^{t}\}$ do
+7:     ${\mathcal{M}}_{c}\leftarrow ConstructExemplarSet(X^{c},n,\theta^{t})$
+8:  end for
+9:  ${\mathcal{M}}\leftarrow({\mathcal{M}}_{1},\dots,{\mathcal{M}}_{N^{t}})$
 4 Empirical evaluation
-We report two settings: (i) Integrated:we evaluate the beneficial contribution of TEAL to competitive ER-IL methods, by replacing their native selection strategy with TEAL (Section 4.2.1); (ii) Alternative selection strategies:we evaluate the beneficial contribution of TEAL as compared to Herding (Section 4.2.2).
-In both cases, we incrementally train a deep-learning model with a fixed-size replay buffer and monitor the average accuracy defined below upon completion of each task $t$.
+We report two settings: (i) Integrated:we evaluate the beneficial contribution of TEAL to competitive ER-IL methods, by replacing their native selection strategy (vanilla version) with TEAL (Section 4.2); (ii) Alternative selection strategies:we evaluate the beneficial contribution of TEAL as compared to other selection strategies (Section 4.2).
+In both cases, we incrementally train a deep model with a fixed-size replay buffer and monitor the average accuracy defined below upon completion of each task $t$.
 4.1 Methodology
-The majority of the experiments are conducted using the open-source CL library Avalanche [16], with the exception of those involving XDER [1], which is not integrated into Avalanche.
+The majority of the experiments are conducted using the open-source Continual Learning library Avalanche (Lomonaco et al.
+2021), with the exception of those involving XDER (Boschini et al.
+2022), which is not integrated into Avalanche.
 For XDER, we utilize the code provided by the authors.
 In order to guarantee a fair comparison and in all experiments, we employ identical network architectures and maintain consistent experimental conditions.
-We make certain in both settings that the buffer remains class-balanced, namely, each update maintains an equal representation of exemplars across all classes.
-When the buffer size is not evenly divisible by the number of classes, there may be instances where certain classes possess one additional exemplar.
-To maintain this balance, we follow this procedure: Prior to adding new exemplars we calculate $n=\frac{|{\mathcal{M}}|}{N^{t}}$, where $N^{t}$ is the total number of existing and new classes.
-Subsequently, we adjust the number of exemplars from existing classes to accommodate $n$ by removing the least likely points, followed by the addition of $n$ exemplars from each new class.
+We make certain in both settings that the buffer remains class-balanced, namely, each update maintains an equal representation of exemplars across all classes [^2] When the buffer size is not evenly divisible by the number of classes, there may classes with an additional exemplar..
+To maintain balance, we follow this procedure: Prior to adding new exemplars we calculate $n=\frac{|{\mathcal{M}}|}{N^{t}}$, where $N^{t}$ is the total number of existing and new classes.
+Subsequently, we adjust the number of exemplars from existing classes to accommodate $n$ by removing redundant points, followed by the addition of $n$ exemplars from each new class.
 In the second setting, we employ a simple baseline ER-IL model.
@@
 Additionally, we utilize a weighted data loader to ensure a balanced mix of data from both the new classes and the exemplars stored in the buffer in each batch.
-Across experiments, the only variation lies in the selection strategy employed: some experiments utilize random-sampling, others utilize Herding, and the remaining experiments leverage TEAL.
-In some experiments, we replace the native buffer selection strategy of an existing ER-IL method with TEAL or Herding to assess their impact.
+Across experiments, the only variation lies in the selection strategy employed: some experiments utilize one of the selection strategies baselines described below, and the remaining experiments employ TEAL.
 Other than this change in the method’s buffer population mechanism, everything else remains the same.
+ER-IL baselines    The following ER-IL methods are used: XDER (Boschini et al.
+2022), which updates the memory buffer by integrating current and past information; ER-ACE (Caccia et al.
+2021), which applies separate losses for new and past tasks; ER (Chaudhry et al.
+2019a), which explores selection strategies in a simple ER framework; BiC (Wu et al.
+2019); iCaRL (Rebuffi et al.
+2017); GEM (Lopez-Paz and Ranzato 2017); and GDumb (Prabhu, Torr, and Dokania 2020).
+To assess the suitability of these methods under the CIL conditions studied here, we used Split CIFAR-100 and set the buffer size to 300, 500, and 2000.
+Figure 3 presents the results for a buffer size of 300 (see Suppl B for results with larger buffer sizes).
+As clearly illustrated in Fig.
+3, GEM and GDumb perform poorly in this scenario, likely due to their unsuitability for a very small memory buffer.
+Similarly, BiC remains competitive only when the buffer size is 2000.
+While iCarl remains competitive across all buffer sizes, it is inherently non-modular, relying heavily on its native Herding selection strategy and a K-Means classifier.
+Hence we cannot integrate it with TEAL.
+We are therefore left with the methods XDER, ER-ACE, and ER, which are both suitable and competitive for further examination with and without TEAL as the selection mechanism.
+For ER-ACE, which requires populating the buffer with exemplars at the start of the task before training on new classes, we cannot rely on the representation provided by a trained model.
+To address this, we adjust the integration of TEAL by filling the buffer in two stages.
+First, at the beginning of task $t$, we populate the buffer with exemplars from the new classes using a random-sampling selection strategy.
+Second, after completing the training for task $t$, we replace the exemplars from the new classes in the buffer with those selected using TEAL.
+It’s important to note that throughout this process, the buffer maintains a fixed size, ensuring compliance with the class-incremental framework.
+Figure 3: Baseline CIL using Split CIFAR-100 with a small buffer size $|{\mathcal{M}}|=300$. 7 baseline methods are shown, reporting the average accuracy $A_{t}$ as new tasks are learned.
+Selection Strategies baselines    We compare with the following selection strategies (see discussion of related work in the introduction): Random sampling, Herding, Uncertainty, and Centered (closest-to-center) - another common strategy that selects the exemplars closest to the center of all elements in feature space.
+Centered focuses on individual exemplars, unlike Herding, which keeps the mean of the whole constructed set close to the center.
+Metrics    As customary, we employ a score that reflects the accuracy during each stage of the incremental training.
+Let $T$ denote the total number of tasks, and $a_{t,i}$ denote the accuracy of task $i$ at the end of task $t$ ( $i\leq t\leq T$).
+For every task $t={1,\dots,T}$, define its average accuracy $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i}$.
+This metric provides a single value for each incremental step, enabling us to directly compare different methods at each step.
+Since by construction each task has the same number of classes, there is no need for additional weighting terms.
+We do not include the metric of forgetting, which estimates how much the model has forgotten about previous tasks, since it is less suitable for the class-incremental setting: with the addition of new classes, performance inevitably drops across all classes (Masana et al.
+2022).
 Datasets    We use several well known Continual Learning benchmarks.
-(i) Split CIFAR-100[22, 5], a dataset created by splitting CIFAR-100 [13] into 10 tasks, each containing 10 different classes of 32X32 images, with 500 images per class for training and 100 for testing.
-(ii) Split tinyImageNet, a dataset created by splitting tinyImageNet [14] into 10 tasks, each containing 20 different classes of 64X64 images, with 500 images per class for training and 50 for testing.
-(iii) Split CUB-200[3], a dataset created by splitting the CUB-200 [26] high-resolution image classification dataset, consisting of 200 categories of birds, with around 30 images per class for training and 30 for testing.
+(i) Split CIFAR-100(Rebuffi et al.
+2017; Chaudhry et al.
+2019b), a dataset created by splitting CIFAR-100 (Krizhevsky, Hinton et al.
+2009) into 10 tasks, each containing 10 different classes of 32X32 images, with 500 images per class for training and 100 for testing.
+We also use a variation of this dataset by splitting CIFAR-100 into 20 tasks instead of 10, as reported in Suppl B.
+(ii) Split tinyImageNet, a dataset created by splitting tinyImageNet (Le and Yang 2015) into 10 tasks, each containing 20 different classes of 64X64 images, with 500 images per class for training and 50 for testing.
+(iii) Split CUB-200(Chaudhry et al.
+2018b), a dataset created by splitting into 20 tasks the CUB-200 (Wah et al.
+2011) high-resolution image classification dataset, consisting of 200 categories of birds, with around 30 images per class for training and 30 for testing.
 Architectures    In the first setting, except for one experiment, we employ a ResNet-18 model in all our experiments.
 The exception is the experiment involving the training of XDER on the Split CUB-200 dataset.
-Due to computational limitations, we used instead a pre-trained ResNet-50 model [11], and this was maintained under all the relevant conditions.
-In the second setting, we use a smaller version of ResNet-18 [11] as a simple baseline ER-IL model [see 17].
-Metrics    As customary, we employ a score that reflects the accuracy during each stage of the incremental training.
-Let $T$ denote the total number of tasks, and $a_{t,i}$ denote the accuracy of task $i$ at the end of task $t$ ( $i\leq t\leq T$).
-For every task $t={1,\dots,T}$, define it average accuracy $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i}$.
-This metric provides a single value for each incremental step, enabling us to directly compare different methods at each step.
-Since all of the experiments we conduct are task-balanced in the sense that each task has an equal number of classes, there is no need to add a weighting term to the metric.
-Baselines    The following methods are suitable for our investigation as they are designed to address ER-IL: XDER [1] updates the memory buffer by adding current information to past memories and preparing the model for new tasks.
-ER-ACE [2], or ER with Asymmetric Cross-Entropy, uses a separate loss for new and past tasks.
-BiC [28] introduces a bias correction layer to address class imbalance in incremental learning.
-iCaRL [22] combines the incremental classifier and representation learning with a nearest-mean-of-exemplars strategy, and Herding for exemplar selection.
-Since the classifier is based on the nearest-mean-of-exemplars, Herding is crucial for the selection strategy, as it guarantees that the features of the exemplars in the buffer approximate the original classes means.
-GEM [17] uses Gradient Episodic Memory to prevent gradient leakage between new and past tasks through constrained optimization.
-GDumb [21] employs a greedy sampling strategy for memory storage and retrains the neural network from scratch during inference.
-ER [4] explores different selection strategies in a simple ER framework.
-In order to evaluate the competitiveness of these methods, we conducted multiple runs on Split CIFAR-100, utilizing fixed buffer sizes of 300, 500, and 2000. Fig. 2 shows the results.
-(a) $|{\mathcal{M}}|=300$
-(b) $|{\mathcal{M}}|=500$
-(c) $|{\mathcal{M}}|=2000$
-Figure 2: Baseline CIL methods comparison on Split CIFAR-100 with different fixed buffer sizes.
-For each method, each point represents the average accuracy achieved on the corresponding task $t$, $A_{t}$.
-Probably because the memory buffer is so small, and as can be seen in Fig.
-2, GEM and GDumb perform poorly in a relative comparison.
-BiC performs well on buffer size 2000, but significantly degrades for smaller buffers for which it is not suited.
-iCarl demonstrates high performance, but it is highly non-modular, being heavily dependent on the Herding selection strategy while using a K-Means classifier.
-Hence we cannot integrate it with TEAL.
-We are therefore left with methods XDER, ER-ACE and ER, which are both suitable and competitive, to be examined with and without TEAL as the selection mechanism.
-For ER-ACE, since it depends on populating the buffer with exemplars at the beginning of the task before training on the new classes, we cannot rely on the representation provided by a trained model.
-Therefore, we adjust the integration of TEAL by filling the buffer in two stages: First, at the beginning of task $t$, we update the buffer with exemplars from the new classes using the random-sampling selection strategy.
-Second, after training task $t$, we replace the exemplars from the new classes in the buffer with exemplars chosen using TEAL.
-It’s important to note that throughout this process, the buffer maintains the fixed size of exemplars it allowed, ensuring compliance with the class-incremental framework.
+Due to computational limitations, we used instead a pre-trained ResNet-50 model (He et al.
+2016), and this was maintained under all the relevant conditions.
+Additionally, we run some experiments with another architecture called ArchCraft, a ResNet variant designed for improved performance in CL (Lu et al.
+2024).
+In the second setting, we use a smaller version of ResNet-18 (He et al.
+2016) as a simple baseline ER-IL model (see Lopez-Paz and Ranzato 2017).
 4.2 Main results
-(b) $|{\mathcal{M}}|=500$
-(c) $|{\mathcal{M}}|=2000$
-Split CIFAR-100
-(a) $|{\mathcal{M}}|=300$
-(e) $|{\mathcal{M}}|=1000$
-(f) $|{\mathcal{M}}|=2000$
-Split tinyImageNet
-(d) $|{\mathcal{M}}|=400$
-(h) ER
-(i) XDER
-Split CUB-200
-(g) ER-ACE
+(a) Split CIFAR-100, $|{\mathcal{M}}|=300$
+(b) Split tinyImageNet, $|{\mathcal{M}}|=400$
+(c) Split CUB-200
 Figure 4: The performance of ER-IL methods with and without TEAL.
-Each plot displays the average accuracy after training incrementally on a different number of classes.
+First row displays the average accuracy after training incrementally on a different number of classes.
 Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy.
 The error bars correspond to standard error based on 4-10 repetitions.
-4.2.1 TEAL integrated with SOTA ER-IL methods
-Split CIFAR-100
-(a) $|{\mathcal{M}}|=300$
-(b) $|{\mathcal{M}}|=500$
-(c) $|{\mathcal{M}}|=2000$
-(e) $|{\mathcal{M}}|=1000$
-(f) $|{\mathcal{M}}|=2000$
-Split tinyImageNet
-(d) $|{\mathcal{M}}|=400$
-(h) ER
-(i) XDER
-Split CUB-200
-(g) ER-ACE
-Figure 4: Performance improvement of TEAL when integrated with different ER-IL methods.
-Each graph depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks.
-The first two rows present the three methods improvement on Split CIFAR-100 and Split tinyImageNet with different fixed buffer sizes denoted as $|{\mathcal{M}}|$.
-The last row pertains to Split CUB-200, where each graph showcases different buffer sizes for a different method.
-Table 1: Final average accuracy for method with and without TEAL
-Dataset Split CIFAR-100 Split tinyImageNet Split CUB-200
-$|{\mathcal{M}}|$ 300 500 2000 400 1000 2000 200 400
-XDER 41.97±0.25 47.97±0.22 57.69±0.21 20.49±0.19 32.68±0.24 39.76±0.31 43.63±1.78
-XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17 23.61±0.32 34.28±0.24 40.76±0.23 49.55±0.47
-Improvement 3.08 2.32 1.1 3.13 1.59 1.0 5.96
-ER-ACE 35.63±0.36 40.6±0.34 53.6±0.17 13.02±0.2 17.37±0.34 21.71(±0.17) 8.56±0.15 11.25±0.99
-ER-ACE+TEAL 40.59±0.2 44.25±0.35 56.11±0.19 14.43±0.29 18.88±0.21 24.12±0.24 10.41±0.16 12.33±0.26
-Improvement 4.97 3.65 2.51 1.41 1.51 2.4 1.85 1.08
-ER 14.82±0.21 18.21±0.29 36.54±0.31 7.75±0.06 7.7±0.06 8.27±0.08 8.98±0.18 12.01±0.52
-ER+TEAL 17.97±0.25 22.39±0.31 38.17±0.37 7.86±0.05 7.99±0.06 8.82±0.09 10.37±0.23 14.13.76±0.3
-Improvement 3.15 4.18 1.62 0.11 0.29 0.56 1.39 2.12
-As mentioned above, we investigate the baseline methods XDER, ER-ACE and ER, comparing their performance when using either their native selection strategies or TEAL.
-We examined Split CIFAR-100 with buffer sizes of 300, 500 and 2000, Split tinyImageNet with buffer sizes of 400, 1000 and 2000, and Split CUB-200 with buffer sizes of 200222Due to computational limitation of the relevant package, we did not run XDER with buffer size 200.
+Second row depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks.
+TEAL integrated with SOTA ER-IL methods
+Table 1: Final average accuracy for method with and without TEAL.
+XDER ER-ACE ER
+Dataset $|{\mathcal{M}}|$ Vanilla +TEAL Improvement Vanilla +TEAL Improvement Vanilla +TEAL Improvement
+Split CIFAR-100 100 27.98±0.18 31.08±0.29 3.1 24.37±0.42 31.06±0.27 6.9 10.31±0.03 10.24±0.03 -0.07
+300 41.97±0.25 45.05±0.24 3.08 35.37±0.29 41.28±0.2 5.92 14.82±0.21 17.97±0.25 3.15
+500 47.97±0.22 50.29±0.2 2.32 40.7±0.25 45.99±0.33 5.29 18.21±0.29 22.39±0.31 4.18
+1000 53.69±0.32 55.02±0.34 1.33 48.16±0.14 51.85±0.22 3.69 26.66±0.81 28.58±0.65 1.92
+2000 57.69±0.21 58.79±0.17 1.1 55.99±0.1 58.7±0.17 2.71 36.54±0.31 38.17±0.37 1.62
+3000 58.96±0.25 60.09±0.25 1.13 59.75±0.25 62.3±0.19 2.55 43.72±1.63 46.03±1.49 2.32
+4000 59.89±0.16 60.42±0.28 0.54 62.6±0.22 64.81±0.2 2.22 49.03±0.89 52.86±0.62 3.82
+Split tinyImageNet 200 14.6±0.16 16.76±0.3 2.16 11.3±0.08 12.47±0.22 1.17 7.95±0.02 8.01±0.03 0.07
+400 20.42±0.12 23.58±0.18 3.16 13.02±0.2 14.43±0.29 1.41 7.75±0.06 7.86±0.05 0.11
+600 25.74±0.09 28.71±0.23 2.97 14.48±0.13 15.7±0.14 1.22 7.54±0.02 7.8±0.07 0.26
+1000 32.68±0.24 34.28±0.24 1.59 17.37±0.34 18.88±0.21 1.51 7.7±0.06 7.99±0.06 0.29
+2000 39.76±0.31 40.76±0.23 1.0 21.71±0.17 24.12±0.24 2.4 8.27±0.08 8.82±0.09 0.56
+4000 43.91±0.14 44.42±0.26 0.51 27.15±0.13 29.62±0.18 2.47 11.61±0.1 13.59±0.21 1.99
+6000 44.71±0.15 45.44±0.39 0.73 30.3±0.23 33.3±0.23 3 16.53±0.15 19.07±0.18 2.54
+Split CUB-200 200 8.56±0.15 10.41±0.16 1.85 8.98±0.18 10.37±0.23 1.39
+400 43.63±1.78 49.55±0.47 5.96 11.25±0.99 12.33±0.26 1.08 12.01±0.52 14.13±0.3 2.12
+As mentioned above, we investigate the baseline methods XDER, ER-ACE, and ER, comparing their performance using either their native selection strategies or TEAL.
+We conducted experiments on Split CIFAR-100 with buffer sizes ranging from 100 to 4000, Split tinyImageNet with buffer sizes from 200 to 6000, and Split CUB-200 with buffer sizes of 200333Due to computational limitations of the relevant package, we did not run XDER with a buffer size of 200.
 and 400.
-Fig.
-3 shows the average accuracy $A_{t}$ after each task $t$.
-Fig.
-4 shows the corresponding improvement of TEAL.
-Finally, the results of the final average accuracy $A_{t}$ and the differences ’improved-vanilla’ are reported in Table 1.
+The results for final average accuracy $A_{t}$ and the differences between ’improved’ and ’vanilla’ are presented in Table 1.
+First row of Fig.
+4 shows the average accuracy $A_{t}$ after each task $t$, while the second row illustrates the corresponding improvement from TEAL in selected experiments.
 As demonstrated in Fig.
@@
 Reassuringly, for the majority of the experiments, the improvement becomes apparent starting from the second task.
-Table 1 shows that the difference (’improvement’ row) is always positive, namely, TEAL always improves the outcome.
-Looking at the improvement row for the XDER method, it is evident that the improvement provided by TEAL increases as the buffer size decreases.
-This trend is also demonstrated for ER-ACE and partially for ER on Split CIFAR-100.
-4.2.2 Comparison of selection strategy
-We begin with a simple ER-IL model on Split CIFAR-100333The order of the classes and the partition to tasks is randomly selected and fixed for all comparisons.
-with various buffer sizes, and three different selection strategies: random, Herding and TEAL.
+Table 1 shows that the difference (indicated in the ’improvement’ row) is consistently positive, demonstrating that TEAL always enhances performance.
+However, it is important to note that when catastrophic forgetting in the vanilla version is too severe, the integration of TEAL can no longer mend the damage.
+Consequently, the enhancements to ER and ER-ACE on the Split tinyImageNet dataset are mostly notable with larger buffer sizes.
+Note that the improvement provided by TEAL tends to increase as the buffer size decreases.
+To further explore the relationship between buffer size and relative improvement, we present a graph showing the final average accuracy improvement across various buffer sizes.
+The results for Split CIFAR-100 and Split tinyImageNet using XDER are illustrated in Fig.
+5 (see results with ER-ACE and ER in Suppl B).
+Clearly, while improvements do occur across all buffer sizes, larger improvements are seen for smaller buffers.
+Figure 5: Performance improvement of TEAL when integrated with XDER over various buffer sizes.
+Comparison of selection strategy
+We begin with a simple ER-IL model on Split CIFAR-100444The order of the classes and the partition to tasks is randomly selected and fixed in all conditions.
+with various buffer sizes, and 5 different selection strategies: random, Herding, Uncertainty, Centered and TEAL.
 As depicted in Fig.
-5, our method enhances performance as compared to random selection by almost 3% for smaller buffers (300, 400, 500) and about 2.5% for larger buffers (1000, 2000).
-When compared to Herding, our method is more effective for smaller buffer sizes, enhancing performance by up to 1.2% for a buffer size of 300, but less effective with large buffers.
-Similar results, using different class ordering, are shown in App A.1.
-Figure 5: Split CIFAR-100: performance difference between TEAL and either random-sampling or Herding selection. Each bar corresponds to a fixed buffer size.
-Dataset Split CIFAR-100
-$|{\mathcal{M}}|$ 300 500 2000
-XDER+TEAL_OneTime 44.13±0.26 49.87±0.22 59.05±0.14
-XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17
-ER-ACE+TEAL_OneTime 35.81±0.42 40.16±0.3 51.52±0.24
-ER-ACE+TEAL 36.4±0.12 40.26±0.23 51.68±0.35
-ER+TEAL_OneTime 17.64±0.12 22.01±0.18 40.65±0.32
-ER+TEAL 18.1±0.1 22.07±0.13 39.96±0.23
-Table 2: The improvement of 2 variants of TEAL when integrated with different CIL methods: XDER, ER-ACE and ER. In almost all cases, the original TEAL matches or outperforms the OneTime variant.
-Given the results shown in Fig.
-5, wherein TEAL seems less effective than herding with large buffers when training the simple baseline model, we repeat this comparison while integrating both methods - TEAL and Herding - into competitive ER-IL methods.
-Fig.
-6 shows the difference between the improvement obtained by TEAL and the one obtained by Herding.
-Since in almost all cases this difference is positive, we may conclude that TEAL is significantly more beneficial than Herding.
+7, our method enhances performance as compared to Uncertainty by 3-4%, and random selection by almost 3% for smaller buffers (300, 400, 500) and about 2.5% for larger buffers (1000, 2000).
+When compared to Herding, our method is more effective with smaller buffer sizes, enhancing performance by up to 1.2% for a buffer size of 300, but less effective with large buffers.
+It is possible that the mechanism of retaining a set of the nearest neighbors around the average sample in each class becomes less effective when the buffer is too small.
+Additionally, Herding lacks a component for selecting a diverse set, which may lead to the retention of similar exemplars, an issue that becomes more pronounced with smaller buffer sizes.
+Interestingly, despite its similarity to Herding, Centered performs worse on larger buffer sizes, resulting in a 1.5-2.5% improvement by TEAL.
+Similar results, using different class ordering, are shown in Suppl A.2.
 (a) $|{\mathcal{M}}|=300$
@@
 Figure 6: The difference between the improvement obtained by TEAL and the one obtained by Herding, showing 2 ER-IL methods and 3 buffer sizes while training on the Split CIFAR-100 dataset.
+Figure 7: Split CIFAR-100: TEAL performance gain compared to 4 baselines: random-sampling, Herding, Centered, and Uncertainty, each bar corresponds to a fixed buffer size.
+The error bars correspond to standard error based on 10-20 repetitions.
+Given the results shown in Fig.
+7, we repeat this comparison while integrating both methods - TEAL and Herding - into competitive ER-IL methods.
+Fig.
+6 shows the difference between the improvement obtained by TEAL and the one obtained by Herding.
+Since in almost all cases this difference is positive, we may conclude that TEAL is significantly more beneficial than Herding.
+Results of Task-IL
+The incorporation of TEAL into existing ER-IL method is also beneficial in the task incremental scenario, although the advantage is less pronounced, likely because the scenario is easier.
+Results are shown in Table 2.
+Table 2: Task-IL, accuracy with and without TEAL
+Dataset $|{\mathcal{M}}|$ XDER XDER+TEAL Improvement
+Split CIFAR-100 100 73.75±0.59 76.56±0.27 2.81
+300 83.5±0.06 84.46±0.26 0.96
+500 85.96±0.14 86.1±0.2 0.14
+Split tinyImageNet 200 42.48±0.38 46.55±0.37 4.07
+400 55.03±0.15 59.2±0.47 4.17
+600 63.69±0.21 65.26±0.28 1.57
+1000 69.72±0.31 70.61±0.23 0.89
 4.3 Ablation study
-We explore an alternative variant of TEAL, which selects the set of exemplars in a single pass.
+Iterative process of TEAL
+We explore an alternative variant of TEAL, which selects the set of exemplars in a single pass instead of by iterations.
 We call this variant TEAL_OneTime.
 While still selecting exemplars for each class separately, this variant first partitions the exemplars of each class into $n=\frac{|{\mathcal{M}}|}{N^{t}}$ clusters, and then selects the most typical exemplar from each cluster in descending order of cluster size, preserving the order of selections.
-The experiments on the two variants are conducted using Split CIFAR-100 with the smaller version of ResNet-18 mentioned in Section 4.1 and buffer sizes of 300, 500 and 2000.
+Table 3: The improvement of 2 variants of TEAL when integrated with different CIL methods: XDER, ER-ACE, and ER. In almost all cases, the original TEAL matches or outperforms the OneTime variant.
+$|{\mathcal{M}}|$ Selection Strategy XDER ER-ACE ER
+300 TEAL_OneTime 44.13±0.26 35.81±0.42 17.64±0.12
+TEAL 45.05±0.24 36.4±0.12 18.1±0.1
+500 TEAL_OneTime 49.87±0.22 40.16±0.3 22.01±0.18
+TEAL 50.29±0.2 40.26±0.23 22.07±0.13
+2000 TEAL_OneTime 59.05±0.14 51.52±0.24 40.65±0.32
+TEAL 58.79±0.17 51.68±0.35 39.96±0.23
+Table 4: Final average accuracy for method with and without TEAL using the ArchCraft architecture. Top 3 rows correspond to Split CIFAR-100, bottom 3 to Split tinyImageNet.
+ER-ACE ER
+$|{\mathcal{M}}|$ Vanilla +TEAL Imp. Vanilla +TEAL Imp.
+300 36.75±0.28 42.47±0.08 5.72 14.59±0.08 18.01±0.35 3.42
+500 41.98±0.24 47.53±0.35 5.56 18.54±0.4 22.3±0.4 3.76
+2000 57.33±0.18 60.18±0.17 2.85 41.03±0.37 43.44±0.28 2.41
+400 12.24±0.11 13.2±0.38 0.96 7.67±0.13 7.84±0.19 0.17
+1000 15.64±0.37 16.7±0.2 1.07 8.12±0.05 8.36±0.18 0.24
+2000 19.08±0.36 20.43±0.37 1.35 9.65±0.11 10.66±0.03 1.01
+The experiments on the two variants are conducted using Split CIFAR-100 with the smaller version of ResNet-18 mentioned in Section 3 and buffer sizes of 300, 500 and 2000.
 We investigate the performance enhancement of each variant when integrated with XDER, ER-ACE, and ER.
 The results are shown in Table 3.
-Whenever there is a significant difference in performance, the iterative variant outperforms the other one in 4 out of 5 cases.
-However, the differences are not substantial.
+Clearly, whenever there is a significant difference in performance, the iterative variant outperforms the other one in 4 out of 5 cases.
+Different architecture
+Some experiments were replicated using ArchCraft instead of the ResNet-18 architecture, obtaining similar results (see Table 4).
 5 Summary and discussion
-We propose a new mechanism to select exemplars for the memory buffer in replay-based CIL methods.
+We proposed a new mechanism to select exemplars for the memory buffer in replay-based CIL methods.
 This method is based on the principles of diversity and representativeness.
-When the memory buffer is relatively small, our method TEAL is shown to outperform both the native mechanism of each ER-IL method (usually random selection) and an alternative selection mechanism called Herding.
-Even when the buffer is large, our method is beneficial in almost all studied cases.
-Our future work will investigate ways to determine at which point, if any, the added value of our method diminishes or possibly becomes harmful.
-We will also integrate TEAL with other replay-based CIL methods that utilize different selection strategies, and analyze the impact on their performance.
-Acknowledgement
-This work was supported by grants from the Israeli Council of Higher Education and the Gatsby Charitable Foundations.
-Appendix
+When the memory buffer is relatively small, our method TEAL is shown to outperform both the native mechanism of each ER-IL method (usually random selection) and alternative selection mechanisms, including Herding, Uncertainty and Centered.
+Even when the buffer is large, our method is beneficial in almost all cases.
+In future work we will investigate ways to determine at which point, if any, the added value of our method diminishes or possibly becomes harmful.
+We will also integrate TEAL with other replay-based CIL methods that utilize different selection strategies.
+Supplementary Material
 Appendix A Implementation details
-The source code used in this study would be uploaded to our GitHub repository and will be made public upon acceptance.
+The source code for this study can be found in the zip file included in the supplementary material.
+This zip file contains a README.md file that provides instructions on how to run the experiments.
+The code would be uploaded to our GitHub repository and will be made public upon acceptance.
+A.1 TEAL implementation
+Clustering algorithm
+We used scikit-learn KMeans implementation.
+Iterations pace
 TEAL, as described in Alg.
@@
 We set a base $b=1.4$, and define $s_{1}=\lfloor b^{4}\rfloor,s_{2}=\lfloor b^{5}\rfloor,\dots,s_{k}=\lfloor b^{log_{b}n}\rfloor$.
-A.1 Stand-alone setting
+A.2 Stand-alone setting
 For all selection strategies, we use a smaller ResNet-18, as mentioned above, trained for 200 epochs.
@@
 We run this setting on Split CIFAR-100 with three different random orders of classes:
-  1. 1.
+1.
 The order we display on Fig.
 7 is: [44, 19, 93, 90, 71, 69, 37, 95, 53, 91, 81, 42, 80, 85, 74, 56, 76, 63, 82, 40, 26, 92, 57, 10, 16, 66, 89, 41, 97, 8, 31, 24, 35, 30, 65, 7, 98, 23, 20, 29, 78, 61, 94, 15, 4, 52, 59, 5, 54, 46, 3, 28, 2, 70, 6, 60, 49, 68, 55, 72, 79, 77, 45, 1, 32, 34, 11, 0, 22, 12, 87, 50, 25, 47, 36, 96, 9, 83, 62, 84, 18, 17, 75, 67, 13, 48, 39, 21, 64, 88, 38, 27, 14, 73, 33, 58, 86, 43, 99, 51]
-  2. 2.
-    The order in 7(a) is: [45, 15, 90, 32, 35, 63, 17, 72, 79, 96, 48, 36, 16, 11, 23, 80, 22, 58, 3, 62, 50, 33, 66, 99, 43, 76, 7, 57, 81, 82, 6, 10, 24, 52, 95, 73, 91, 21, 38, 31, 85, 59, 13, 69, 75, 70, 64, 8, 77, 34, 46, 39, 92, 0, 44, 98, 49, 9, 4, 61, 12, 83, 28, 78, 40, 88, 54, 5, 26, 41, 89, 20, 84, 2, 1, 55, 19, 74, 25, 37, 42, 14, 30, 18, 67, 71, 68, 27, 60, 51, 29, 56, 93, 47, 97, 94, 86, 87, 65, 53]
-  3. 3.
-    The order in 7(b) is: [48, 97, 1, 81, 90, 49, 10, 8, 7, 20, 70, 73, 75, 14, 91, 38, 47, 21, 74, 52, 80, 98, 59, 12, 71, 85, 6, 34, 55, 82, 95, 63, 78, 15, 94, 60, 99, 76, 25, 40, 88, 0, 62, 96, 87, 51, 16, 18, 9, 19, 29, 45, 86, 53, 56, 31, 28, 61, 30, 33, 4, 67, 64, 58, 50, 54, 3, 13, 37, 27, 66, 77, 84, 69, 2, 41, 22, 92, 42, 44, 11, 36, 46, 79, 65, 72, 23, 17, 39, 5, 89, 35, 24, 83, 43, 57, 93, 32, 68, 26]
+2.
+The order in Fig.
+8(a) is: [45, 15, 90, 32, 35, 63, 17, 72, 79, 96, 48, 36, 16, 11, 23, 80, 22, 58, 3, 62, 50, 33, 66, 99, 43, 76, 7, 57, 81, 82, 6, 10, 24, 52, 95, 73, 91, 21, 38, 31, 85, 59, 13, 69, 75, 70, 64, 8, 77, 34, 46, 39, 92, 0, 44, 98, 49, 9, 4, 61, 12, 83, 28, 78, 40, 88, 54, 5, 26, 41, 89, 20, 84, 2, 1, 55, 19, 74, 25, 37, 42, 14, 30, 18, 67, 71, 68, 27, 60, 51, 29, 56, 93, 47, 97, 94, 86, 87, 65, 53]
+3.
+The order in Fig.
+8(b) is: [48, 97, 1, 81, 90, 49, 10, 8, 7, 20, 70, 73, 75, 14, 91, 38, 47, 21, 74, 52, 80, 98, 59, 12, 71, 85, 6, 34, 55, 82, 95, 63, 78, 15, 94, 60, 99, 76, 25, 40, 88, 0, 62, 96, 87, 51, 16, 18, 9, 19, 29, 45, 86, 53, 56, 31, 28, 61, 30, 33, 4, 67, 64, 58, 50, 54, 3, 13, 37, 27, 66, 77, 84, 69, 2, 41, 22, 92, 42, 44, 11, 36, 46, 79, 65, 72, 23, 17, 39, 5, 89, 35, 24, 83, 43, 57, 93, 32, 68, 26]
 (a) Order 2
 (b) Order 3
-Figure 7: Enhanced accuracy of TEAL compared to random-sampling and Herding selection strategies, additional two different classes order.
-A.2 Integrated setting
-Here we train a ResNet-18 model for 100 epochs using the same optimizer as in A.1, with the same batch size and data augmentations, with some exceptions.
-ER-ACE starts with a learning rate of 0.01, and all experiments on Split CUB-200 are conducted for 30 epochs with a batch size of 16 due to the dataset size and the resolution of its images.
-A.3 Baselines setting
-The setting for the experiments of the baseline methods is the same as in A.2 with some exceptions.
+Figure 8: Split CIFAR-100: TEAL Enhanced accuracy of TEAL compared to 4 baselines: random-sampling, Herding, Centered, and Uncertainty, in two different classes order.
+A.3 Integrated setting
+Here we train a ResNet-18 model for 100 epochs using the same optimizer as in A.2, with the same batch size and data augmentations, with some exceptions.
+ER-ACE starts with a learning rate of 0.01 and train on batch size of 10 examples as in the original paper, and all experiments on Split CUB-200 are conducted for 30 epochs with a batch size of 16 due to the dataset size and the resolution of its images.
+In experiments using the ArchCraft model, we used the ResAC-A model as implemented in the official paper’s code.
+A.4 Baselines setting
+The setting for the experiments of the baseline methods is the same as in A.3 with some exceptions.
 For BiC, the number of training epochs is 250, and the learning rate scheduler decays the learning rate by a factor of 0.1 on epochs 100, 150 and 200.
@@
 For iCaRL the learning rate starts from 2, the weight decay is 0.00001 and the learning rate scheduler decays the learning rate by a factor of 0.2 on epochs 49 and 63.
-Appendix B Compute resources
+A.5 Compute resources
 All experiments involved training deep learning models, necessitating the use of GPUs.
@@
 Any other experiments conducted for the full research and not reported in the paper required the same compute resources.
+Appendix B Additional Results
+(a) Split CIFAR-100
+(b) Split tinyImageNet
+Figure 9: Performance improvement of TEAL when integrated with ER and with ER-ACE over various buffer sizes.
+(a) Split CIFAR-100, $|{\mathcal{M}}|=500$
+(b) Split CIFAR-100, $|{\mathcal{M}}|=2000$
+(c) Split tinyImageNet, $|{\mathcal{M}}|=1000$
+(d) Split tinyImageNet, $|{\mathcal{M}}|=2000$
+(e) Split CUB-200, $|{\mathcal{M}}|=400$
+Figure 10: The performance of ER-IL methods with and without TEAL.
+First row displays the average accuracy after training incrementally on a different number of classes.
+Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy.
+The error bars correspond to standard error based on 4-10 repetitions.
+Second row depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks.
+(a) $|{\mathcal{M}}|=300$
+(b) $|{\mathcal{M}}|=500$
+(c) $|{\mathcal{M}}|=2000$
+Figure 11: Split CIFAR-100: The performance of ER-IL methods with and without TEAL with splitting CIFAR-100 into 20 tasks instead of 10.
+Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy.
+The error bars correspond to standard error based on 6 repetitions.
+(a) $|{\mathcal{M}}|=500$
+(b) $|{\mathcal{M}}|=2000$
+Figure 12: Baseline CIL methods comparison on Split CIFAR-100 with different fixed buffer sizes.
+For each method, each point represents the average accuracy achieved on the corresponding task $t$, $A_{t}$.